please help me solve this equation for x and y:
3x+7=2y
-2y-1=10x

Answer by dukesmart(20)

3x+7=2y ===========> 2y= 3x+7 ......(1)
-2y-1=10x========== -2y= 10x+1 .....(2)
Add (1) to (2) we get:
2y-2y=3x+10x+7+1 ========> 13x=-8
x=-8/13
y= -(10x+1)/2 = -[-80/13 +1]/2= 67/6